Nitrogen has an atomic mass of 14.0067, its naturally occuring isotopes are 14N (14.003) and 15N (15.001). What are the relative abundance of each isotope?
1
Expert's answer
2019-10-28T07:50:27-0400

Nitrogen has an atomic mass of 14.0067 u.

Let one of its isotope having atomic mass 14.003 u has natural abundance of x %.

Other isotope having atomic mass 15.001 u has natural abundance of (100-x) %.

"(x\u00d714.003 + (100-x)\u00d7 15.001)\/100 =14.0067"

"14.003x +1500.1 -15.001x = 1400.67"

"0.998x=99.43"

"x = 99.43\/0.998"

"x = 99.629"

"100-x =0.371"

Relative abundance of isotope having atomic mass 14.003 is 99.629 %(Ans)

The other isotope has abundance of 0.371 %(ans)